OpenStudy (anonymous):

Trent purchases 44 euros worth of souvenirs while on vacation in France. If $1 U.S. = 0.678 euros , find the cost of the souvenirs in US dollars. Round to the nearest cent

OpenStudy (texaschic101):

1/.678 = x/44 ($1 to .678 euro = x$ to 44 euros) cross multiply .678x = 44 x = 44/.678 x = 64.896 rounds to 64.90

OpenStudy (mathstudent55):

Here is another approach at a conversion of units. Since $1 = 0.678 E, then by dividing both sides by 0.678 E you get \( \dfrac{$1}{0.678 ~E} = 1\) Since multiplication by 1 does not change a number, then \(44 E \times \dfrac{$1}{0.678 ~E} \) is still equal to 44 E, but once we multiply it out, the units change giving us US$.